### v2

# Transferable Mask Transformer: Cross-domain Semantic Segmentation with Region-adaptive Transferability Estimation

Abstract
Recent advances in Vision Transformers (ViTs) have set new benchmarks in semantic segmentation. However, when adapting pretrained ViTs to new target domains, significant performance degradation often occurs due to distribution shifts, resulting in suboptimal global attention. Since self-attention mechanisms are inherently data-driven, they may fail to effectively attend to key objects when source and target domains exhibit differences in texture, scale, or object co-occurrence patterns. While global and patch-level domain adaptation methods provide partial solutions, region-level adaptation with dynamically shaped regions is crucial due to spatial heterogeneity in transferability across different areas of the image. In this paper, we present Transferable Mask Transformer (TMT), a novel region-level adaptation framework for semantic segmentation, which aligns cross-domain representations through spatial transferability analysis. TMT consists of two key components: (1) An Adaptive Cluster-based Transferability Estimator (ACTE), which dynamically segments images into structurally and semantically coherent regions for localized transferability assessment, and (2) A Transferable Masked Attention (TMA) module, which integrates region-specific transferability maps into ViTs’ attention mechanisms, prioritizing adaptation in regions with low transferability and high semantic uncertainty. Comprehensive evaluations across 20 cross-domain pairs demonstrate TMT’s superiority, achieving an average $2\%$ MIoU improvement over vanilla fine-tuning and a $1.28\%$ increase compared to the state-of-the-art baselines. The source code will be publicly available soon.
1 Introduction
Refer to caption Refer to caption Refer to caption
(a) Examples of different datasets
Vanilla Finetuning
pOursp
(b) Attention Mask
Vanilla Finetuning
pOursp
(c) Attention Map
Equal-sized Patch
Adaptive Region
(d) Partition Method
Figure 1: Illustration of Different Datasets and Attention Mechanisms: Fig.a shows the domain gap across different datasets, highlighting the challenges encountered during data collection in various real-world scenarios. In Fig.b-c, the attention mask and attention map visualizations indicate that vanilla fine-tuning models tend to predict misleading masks, focusing on less relevant regions (highlighted in red). In contrast, our method generates more accurate masks (shown in blue), directing attention to more critical regions in the target task, such as the car. Fig.d compares two region partition methods for estimating transferability, where our approach segments the scene into adaptive regions.
Semantic segmentation, a core task in computer vision, assigns semantic categories to each pixel in an image. While vision transformers [10, 29] have surpassed CNN-based models and achieved state-of-the-art performance, fine-tuning these models remains challenging due to distribution shifts[16, 25, 14, 9, 19]. As shown in Fig. 1(a), domain gaps arise when source data differs significantly from the target domain due to variations in visual features like lighting, background, or object appearances. These shifts are particularly problematic for vision transformers, whose global attention mechanisms can be easily disrupted. For instance, Fig. 1(b) and 1(c) demonstrates that vanilla fine-tuning often leads to misdirected attention maps, where the model fails to attend to semantically related patches. This phenomenon aligns with previous observations [26] that distribution shifts can degrade model performance due to difference in object co-occurrence patterns. Therefore, developing a more fine-grained fine-tuning method is essential to address these challenges and achieve robust performance on unseen and domain-diverse segmentation tasks.
To address the challenges of distribution shifts in vision transformers, several approaches have been proposed. For instance, methods like [26] and [16] focus on aligning feature distributions between source and target domains, while others [9, 19] employ more advanced fine-tuning strategies to mitigate domain gaps. However, these methods often treat the entire image as a uniform unit, not taking into account the spatial heterogeneity in transferability across regions in images from different domains [28, 30, 4]. For example, in autonomous driving scenarios, background regions like the sky are highly consistent across domains and thus easily transferable, while intricate urban elements pose greater challenges with lower transferability due to their complexity and variability. This discrepancy suggests that a global or patch-level approach to fine-tuning is insufficient, as it overlooks the need for region-adaptive transferability assessment.
To address this limitation, we propose a novel Adaptive Cluster-based Transferability Estimator (ACTE), which flexibly evaluates region-level transferability. While global and patch-level domain adaptation methods offer partial solutions [28, 31, 23], they often fall short in capturing regional semantics and tend to disrupt structural and semantic consistency, especially at object boundaries or for small objects, as demonstrated in Fig. 1(d). In contrast, our method introduces a hierarchical and adaptive approach that dynamically segments images into regions based on semantic and structural coherence. By preserving object boundaries and handling small objects effectively, ACTE ensures a localized assessment of transferability, therefore significantly improving cross-domain generalization.
While existing methods for leveraging transferability have demonstrated success in CNN-based architectures [23], their direct application to vision transformers remains underexplored. To bridge this gap, we propose Transferable Masked Attention (TMA), a novel module tailored for transformer-based models. TMA dynamically adjusts attention masks by incorporating region-specific transferability maps, enabling object queries to focus on regions with both low transferability and low semantic confidence. By combining transferability scores in a hierarchical way, TMA ensures that the model prioritizes regions requiring adaptation while suppressing attention to well-aligned or confidently predicted areas. This approach not only mitigates inappropriate mask predictions but also enhances the model’s ability to capture domain-specific details, offering a principled way to integrate transferability estimation into the attention mechanism of vision transformers.
Our experimental results show that the proposed methods achieve an average MIoU improvement of 2% compared to vanilla fine-tuning and 1.28% over the SOTA method across 20 source-target transfer pairs over 5 popular semantic segmentation benchmarks for autonomous driving.
Figure 2: Overview of the framework. The model training begins with the ACTE, which is first trained using both source and target data. The lower section illustrates how ACTE evaluates and assigns different transferability scores to various regions, represented by different colors in the transferability map. After ACTE has been trained, its output—region-level transferability maps—is used to guide the training of the TMA within attention mechanism (top-right).
2 Related Work
2.1 Transformer-Based Visual Segmentation
Visual segmentation is a fundamental problem in computer vision and involves numerous real-world applications, such as robotics, social media, autonomous driving, etc. [18]. Recently, researchers applied transformers to Computer Vision (CV) tasks. Vision Transforemer (ViT) directly takes the sequence of image patches to classify the full image [7]. It achieves state-of-the-art performance on multiple image recognition datasets. DETR [2] simplified object detection with object queries. Mask2Former [5] further improved training efficiency by introducing mask attention, reducing query ambiguity, and limiting cross-attention scope. Despite the success of Mask2Former, challenges arise in domain adaptation, where attention masks derived from the source domain may not generalize well to different target domains due to variations in visual features.
2.2 Transfer Learning in Transformers
Before ViTs, domain-invariant representation learning strategies were used [32]. DAFormer introduced a transformer backbone with rare class sampling, ImageNet feature loss, and learning rate warm-up, becoming a strong baseline [11]. HRDA [12] improved on DAFormer with multi-resolution training, utilizing different crops to maintain fine segmentation details. MIC [13] applied masked image consistency to enhance spatial context relations by enforcing consistency between masked target images and pseudo-labels through a teacher-student framework. SFA [27] and DADETR [34] focused on feature alignment with transformers, employing domain-specific queries and hybrid attention modules. These methods are generally based on unsupervised domain adaptation strategies, aiming to reduce distribution discrepancies between the source and target domains.
2.3 Transferability Estimation in Vision Tasks
Recent advancements in transferability estimation have significantly improved the efficiency and accuracy of transfer learning in vision tasks. OTCE metrics [24] introduced auxiliary-free methods to evaluate transferability with reduced computational costs. Yang et al. extended these concepts to semantic segmentation by using pixel-wise transferability scores, improving fine-tuning accuracy by focusing on challenging low-transferability regions [23]. In object detection, Chen et al. propose HTCN [3] to balance transferability and discriminability, optimizing feature representations at multiple levels. Yang et al. introduced TADA [31], a method that aligns only the most transferable regions within images. Lastly, Yang et al. proposed the Transferable Vision Transformer (TVT), leveraging ViTs to enhance transferability in unsupervised domain adaptation, which differs from ours as we employ supervised fine-tuning.
3 Preliminary
3.1 Supervised Finetuning
For segmentation tasks, the input space $X=\{\bm{x}\colon P\to\mathbb{R}\}$ is the image set, and the output space $Y=\{\bm{y}\colon P\to\Lambda\}$ is the set of partition of images in label, $P=\{1,\dots,h\}\times\{1,\dots,w\}$ is the pixel set. Source domain $\mathcal{D}_{\mathrm{S}}$ and target domain $\mathcal{D}_{\mathrm{T}}$ are different distributions over $X\times Y$. Given a source dataset $\mathcal{S}=\{(\bm{x}_{\mathrm{S}}^{i},\bm{y}_{\mathrm{S}}^{i})\}_{i=1}^{n}$ drawn i.i.d. from $\mathcal{D}_{\mathrm{S}}$ and a target dataset $\mathcal{T}=\{(\bm{x}_{\mathrm{T}}^{i},\bm{y}_{\mathrm{T}}^{i})\}_{i=1}^{n^{\prime}}$ drawn i.i.d. from $\mathcal{D}_{\mathrm{T}}$. We aim to build a classifier $\sigma\colon X\to Y$ with a low target risk
$R_{\mathcal{D}_{\mathrm{T}}}(\sigma):=\operatorname{E}\mathop{\lvert}\{p\in P\mathop{|}\sigma(\bm{x}_{\mathrm{T}})(p)\neq\bm{y}_{\mathrm{T}}(p)\}\mathop{\rvert}$, (1)
in which $(\bm{x}_{\mathrm{S}},\bm{y}_{\mathrm{S}})\sim\mathcal{D}_{\mathrm{S}},\ (\bm{x}_{\mathrm{T}},\bm{y}_{\mathrm{T}})\sim\mathcal{D}_{\mathrm{T}}$. Fine-tuning involves initializing the target model $\theta_{t}$ with the source model parameters $\theta_{s}$, followed by supervised learning on the target data. In the context of supervised finetuning for segmentation tasks, the target risk $R_{\mathcal{D}_{\mathrm{T}}}(\sigma)$ quantifies the expected prediction error of the classifier $\sigma$ on the target domain. Specifically, it measures the average number of misclassified pixels across the target dataset. This risk is typically computed using a pixel-wise loss function, such as the Cross-Entropy Loss or Dice Loss.
3.2 Measuring Transferability with Proxy A-distance
Let $\mathcal{D}_{{}^{X}}$ denote the marginal distribution of $\mathcal{D}$ over $X$. Let $\mathcal{D}_{{}^{X}}$ denote the marginal distribution of $\mathcal{D}$ over $X$. The transferability between domains $\mathcal{D}_{\mathrm{S}}^{X}$ and $\mathcal{D}_{\mathrm{T}}^{X}$ is measured using Proxy $\mathcal{A}$-distance (PAD) [8]. This distance metric quantifies the divergence between source and target domains, providing a principled approach to assess domain shift.
Given $\mathcal{A}$ as a collection of measurable sets on $X$, the $\mathcal{A}$-divergence between $\mathcal{D}_{\mathrm{S}}^{X}$ and $\mathcal{D}_{\mathrm{T}}^{X}$ is given by
$d_{\mathcal{A}}(\mathcal{D}_{\mathrm{S}}^{X},\mathcal{D}_{\mathrm{T}}^{X}):=2\sup_{A\in\mathcal{A}}\left|\operatorname*{P}[\bm{x}_{\mathrm{S}}\in A]-\operatorname*{P}[\bm{x}_{\mathrm{T}}\in A]\right|.$ (2)
We estimate it with the $\mathcal{A}$-divergence between $\mathcal{S}^{X}$ and $\mathcal{T}^{X}$,
$\begin{split}&\!\!\!\!d_{\mathcal{A}}(\mathcal{S}^{X},\mathcal{T}^{X})=\\ &\!\!\!\!2\Bigl{[}1-\min_{A\in\mathcal{A}}\Bigl{(}\frac{1}{n}\!\sum_{1\leqslant i\leqslant n}\!\mathrm{I}[\bm{x}_{\mathrm{S}}^{i}\!\notin\!A]+\frac{1}{n^{\prime}}\!\sum_{1\leqslant i\leqslant n^{\prime}}\!\mathrm{I}[\bm{x}_{\mathrm{T}}^{i}\!\in\!A]\Bigr{)}\Bigr{]},\end{split}$ (3)
in which $\mathcal{S}^{X}=\{\bm{x}\mathop{|}\mathop{\exists}\bm{y}\operatorname{s.t.}(\bm{x},\bm{y})\in\mathcal{S}\}$ for a dataset $\mathcal{S}$. $\mathrm{I}[p]$ is the indicator function. If predicate $p$ is true, $\mathrm{I}[p]=1$, otherwise $\mathrm{I}[p]=0$.
Although computing $d_{\mathcal{A}}(\mathcal{S}^{X},\mathcal{T}^{X})$ exactly is generally challenging, an approximation can be efficiently obtained by training a classifier to distinguish between source and target examples [15]. Define the labeled dataset as
$\mathcal{L}=\{(\bm{x}_{\mathrm{S}}^{i},0)\}_{i=1}^{n}\cup\{(\bm{x}_{\mathrm{T}}^{i},1)\}_{i=1}^{n^{\prime}},$ (4)
where source and target samples are labeled 0 and 1, respectively. This yields the Proxy $\mathcal{A}$-distance as an estimate of $\mathcal{A}$-divergence:
$\hat{d}_{\mathcal{A}}=2(1-2\epsilon).$ (5)
A domain discriminator $E$ trained on $\mathcal{L}$ provides a generalization error $\epsilon$, which approximates the minimization term in Equation (3).
4 Method
Our framework, TMT, illustrated in Fig. 2, builds upon the Mask2Former architecture to address cross-domain image segmentation through a unified and layer-wise adaptive pipeline. The process begins with feature extraction, where a backbone network captures hierarchical features from the input image, followed by a pixel decoder that refines and upsamples these features to preserve spatial details. These feature maps are then processed by the Adaptive Clustering and Transferability Estimator (ACTE), which adaptively segments the image into regions based on structural characteristics and assigns transferability scores to each region. These scores, encoded in a transferability map, quantify domain alignment and guide the subsequent learning process. Crucially, our Transferable Masked Attention (TMA) mechanism integrates this transferability map into every layer of the network, dynamically modulating the attention mechanism to enhance domain adaptation at multiple feature levels. Finally, the refined features and masks are used to generate class predictions and segmentation outputs, enabling accurate and domain-adaptive image segmentation. By incorporating transferability-aware modulation at each layer, our end-to-end pipeline ensures robust performance across domains, explicitly addressing domain shifts at both regional and hierarchical levels.
4.1 Adaptive Cluster Transferability Estimator
Adaptive Clustering Method.  To capture inter-regional structural differences while preserving intra-regional structural and semantic consistency, our approach generates adaptively sized regions that align closely with image boundaries. As illustrated in Fig.3, we propose an adaptive clustering method inspired by the classic works on superpixel learning [1, 17], which significantly reduces the time complexity by confining the computation range of the pixel similarity matrix. We perform an iterative grouping algorithm that takes a feature map $\bm{\chi}$ as input, and gives out a region-level feature map $\bm{f}$, where $R$ denotes the set of regions. We define a soft assignment $\text{\large$a$}\colon P\times R\to[0,1]$, where $\text{\large$a$}(p,r)$ indicates the probability of pixel $p$ being assigned to region $r$.
Figure 3: Pixel features are compared to nearby center features to compute similarities. Each pixel is assigned to one of its nearby center features based on their similarities.
Initially, the pixel space $P$ is divided into cells of size $c\times c$, which serve as clustering centers. To assign coordinates to regions, the region set is defined as $R=\{1,\dots,h/c\}\times\{1,\dots,w/c\}$. Each pixel is assigned to its corresponding cell, resulting in the initial assignment $a$ as follows:
$\text{\large$a$}((x,\!y),(i,\!j))=\mathrm{I}[\text{$(i\!-\!1)c\hskip-0.83328pt<\hskip-0.83328ptx\hskip-0.83328pt\leqslant\hskip-0.83328ptic\text{ and }(j\!-\!1)r\hskip-0.83328pt<\hskip-0.83328pty\hskip-0.83328pt\leqslant\hskip-0.83328ptjr$}].$ (6)
The algorithm then repeats following steps for $L$ iterations. First, the region-level feature map $\bm{f}$ is computed by averaging the features within each region:
$\bm{f}(r)=\frac{1}{\sum_{p\in P}\text{\large$a$}(p,r)}\sum\nolimits_{p\in P}\text{\large$a$}(p,r)\bm{\chi}(p).$ (7)
Next, the similarity between pixels and clustering centers is computed. For computational efficiency, we limit this comparison to the $3\times 3$ grid of nearby centers surrounding each pixel, as illustrated in Fig.3. The similarity function $s\colon P\times R\to[0,+\infty)$ is defined as:
[…]
Here, $\kappa$ is a hyperparameter that controls the sharpness of the similarity distribution. This approach ensures that clustering centers only aggregate nearby pixels, forming cohesive region-level image segments. The new soft assignment is then computed as:
$\text{\large$a$}_{\text{new}}(p,r)=\frac{s(p,r)}{\sum_{\pi\in R}s(p,\pi)}.$ (9)
This updated assignment is substituted back into Equation (7) to recalculate the region-level feature map $\bm{f}$. After $\bm{L}$ iterations, the process converges to a final cluster region feature map. Each position’s feature vector $\bm{f}(r)$ corresponding to a region with a certain receptive field in the original image.
Region-level Transferability Estimator.  Building upon the theoretical framework of Proxy A-distance, we propose a region-level transferability estimation mechanism. To extend this concept to region-level adaptation, we design an domain discriminator $E$ that operates on adaptive regions $\bm{f}(r)$ rather than entire images. Given region-level features $\bm{f}(r)$ extracted from both domains, $E$ predicts the probability $E(\bm{f}(r))\in[0,1]$ that region $r$ originates from the target domain. To calculate the generalization error $\epsilon$ in Equation (5), we can use the cross-entropy loss, defined by:
$\mathcal{L}(E(\bm{f}(r)),d)=(1-d)\log\frac{1}{1\!-\!E(\bm{f}(r))}+d\log\frac{1}{E(\bm{f}(r))},$ (10)
where $d$ represents the domain label of image $\bm{x}\in\mathcal{L}^{X}$ that generates $\bm{f}$, indicating whether $\bm{x}$ is drawn from the source domain ($d=0$) or from the target domain ($d=1$). To train $E$, we construct a training set consisting of features $\bm{F}$ extracted from both the source and target domains, with corresponding domain labels $d$. Training data is constructed by sampling balanced batches from both domains, with features $\bm{f}(r)$ generated by our adaptive cluster module.
A lower value of $E(\bm{f}(r))$ indicates that the features are more similar to the source domain, which is labeled as 0. So we define the transferability map $\bm{T}$ for each pixel as:
$\bm{T}_{x,y}=1-E\left(\bm{f}\left(r^{*}\right)\right),$ (11)
where $r^{*}=\operatorname*{arg\ max}_{r\in R}\text{\large$a$}((x,y),r)$. A high $\bm{T}_{x,y}$ value indicates near-perfect domain alignment, suggesting that the region’s feature distribution closely matches the source domain. Such regions require minimal adaptation, as they already exhibit strong transferability and provide limited room for improvement through fine-tuning. Conversely, regions with lower $\bm{T}_{x,y}$ exhibit significant distribution shifts from the source domain, necessitating substantial and targeted fine-tuning to bridge the domain gap. By focusing on these low-transferability regions, we can effectively adapt the model to capture the unique characteristics of the target domain, thereby enhancing the model’s cross-domain generalization ability.
4.2 Transferability Guided Masked Attention
With the region-level transferability map $\bm{T}$ obtained from the domain discriminator $E$, we now leverage this critical information to guide the transformer’s learning process. The transferability map $\bm{T}$ encodes the degree of domain alignment for each region, providing a principled way to modulate the attention mechanism in vision transformers. Building upon Mask2Former’s architecture [5], we introduce a key innovation: the integration of region-aware transferability maps into the masked attention mechanism, enabling dynamic adaptation to domain shifts.
Formally, the Transferable Masked Attention (TMA) operation is formulated as:
[…]
where $\bm{M}_{i,j}$ and $\bm{T}_{i,j}$ represent the predicted attention mask from the transformer decoder and the transferability scores, respectively. TMA dynamically adjusts the attention mechanism based on both semantic uncertainty and domain divergence, focusing on regions with low semantic confidence ($\bm{M}_{i,j}\leqslant\lambda_{M}$) and low transferability ($\bm{T}_{i,j}\leqslant\lambda_{\mathrm{T}}$). The thresholds $\lambda_{M}$ and $\lambda_{\mathrm{T}}$ control the strictness of this selection process, as analyzed in Section 5.6. This dual-conditioned masking ensures that the model prioritizes regions where adaptation is most critical, while suppressing attention to well-predicted and domain-aligned areas.
  Domain MIoU fwIoU mACC
[…]
Table 1: Performance comparison of different methods across various datasets when transferred from different source domains.
5 Experiments
5.1 Datasets
Cityscapes dataset [6] includes 5000 images of 2048×1024 pixels from 50 German cities, diverse in seasons, times, backgrounds and weather. BDD dataset [33] contains 7000 training and 1000 testing images of 1280×720 pixels, collected from US street scenes. Mapillary [20] is a large-scale dataset with 18000 training images, 2000 validation images, and 5000 testing images from around the world, showcasing high diversity in weather and seasonal conditions. For synthetic datasets, GTAV [21] contains highly realistic images generated from Grand Theft Auto V at 1914×1052 pixels, with 12,403 training, 6,382 validation, and 6,181 testing images. Additionally, SYNTHIA [22] has 9400 images of 1280×760 pixels.
5.2 Training Details
To achieve true transfer accuracy in each transfer experiment, we train the source model for a total of 20,000 iterations, evaluating the test accuracy every 1,000 steps. During the transfer training phase, the estimator is first trained and then frozen to stabilize its evaluation. We then initialize the target model with the pre-trained weights from the source model and proceed to fine-tune the entire model using the AdamW optimizer, with a learning rate of 0.0001 and a batch size of 16. The model training is conducted using two NVIDIA A800 GPUs. The cluster algorithm repeats its steps for 6 iterations.
Figure 4: Visualization of segmentation results where models pretrained on Cityscapes are transferred to the BDD dataset.
5.3 Experimental Results
In our experiments, we utilized five different source datasets—Cityscapes, BDD, Mapillary, SYNTHIA, and GTAV—and evaluated the segmentation performance of models to various target datasets. Each dataset represents different challenges: Cityscapes and BDD are real-world urban datasets, Mapillary offers diverse scenes, SYNTHIA and GTAV are synthetic datasets with varying degrees of realism. The purpose of using such a diverse set of datasets is to assess how well the models can generalize across different domains, particularly from synthetic to real-world scenarios.
Our experimental design incorporates two prevalent adaptation strategies: (1) partial network adaptation through linear probing and (2) complete parameter fine-tuning. These established baselines provide fundamental performance references for domain adaptation tasks. To enhance methodological comparability, we further integrate two state-of-the-art transferability-aware approaches: OTCE-Finetuning and LEEP-Finetuning [24].
5.3.1 Main Results
Our proposed method consistently shows superior performance across the majority of source-target pairs when compared to baseline in Tab.1. The MIoU across the 20 source-target pairs in our experiments shows an average increase of 2% compared to full fine-tuning and an average increase of 1.28% compared to the OTCE-finetuning method.
When Cityscapes or BDD is used as the source, our method consistently achieves the highest scores across all target datasets. This demonstrates the strong generalization ability of our approach when transferring from well-curated, real-world datasets. SYNTHIA and GTAV, being synthetic datasets, both showcase our method’s robust adaptability and mitigate the domain gap, particularly in handling synthetic-to-real domain shifts, which are typically challenging. Mapillary’s results in MIoU are less favorable, likely due to the datasets high complexity and diversity, which can amplify noise.
As shown in Fig.4, our method more accurately delineates the boundaries of vehicles and roads, preserving the complete structure of objects. It also demonstrates a superior ability in segmenting small objects.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Visualization of transferability maps between the Cityscapes (target domain) and Mapillary (source domain) datasets. The left column displays the original input images, while the right column shows the corresponding transferability maps. The maps indicate regions with high transferability in deep red and low transferability in more light color.
5.4 Visualization Analysis
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Vanilla Finetuning Ours Comparison Vanilla Fine-tuning Ours
Figure 6: Visual comparison of attention mechanism: The first two columns show the differences in mask predictions, with the vanilla model (red) broadly dispersing attention, including irrelevant areas like the sky. In contrast, our method (blue) tightens the focus on relevant regions such as the road. The final two columns further illustrate that the vanilla approach misdirects attention to less critical areas, while our method more effectively channels attention towards the essential elements of the task.
5.4.1 Transferability Maps
The transferability maps generated from the experiments clearly demonstrate the effectiveness of our method. Each pair of images in Fig.5 shows the original input on the left and the corresponding transferability map on the right. The target domain is Cityscapes, and the source domain is Mapillary, both of which consist of real urban scenes. Therefore, elements such as vehicles are relatively similar across both domains. This observation is consistent with real-world scenarios, as the features of roads and skies are relatively simple and consistent across different domains, while the features of buildings are more complex and variable. This outcome reflects that the domain discriminator effectively identifies the complex and diverse image information present in urban environments and accurately assesses the transferability of different regions.
5.4.2 Attention Mechanism
We aim to provide a detailed analysis of the internal mechanisms driving the effectiveness of our method in Fig 6. The first column highlights red regions where the vanilla finetuning misdirects masks to less relevant areas, such as the sky, leading to suboptimal segmentation, particularly in tasks like autonomous driving where road layout understanding is critical. In contrast, the second column shows our methods segmentation masks, with blue regions indicating a more accurate focus on important areas. The fourth and fifth columns further demonstrate how our method directs attention to crucial areas, such as the road, rather than irrelevant regions.
  CB CM CS CG BC BM BS BG
 w/o ACTE 63.5 71.7 76.8 70.2 56.5 48.0 54.8 53.8
w/o TMA 63.0 71.2 76.3 69.7 56.0 47.5 54.3 53.3
Ours 64.5 72.7 77.8 71.2 57.5 49.0 55.8 54.8
Table 2: Ablation Study Results.
5.5 Ablation Study
We conducted an ablation study on the transfer from Cityscapes and BDD to other datasets in Tab.2. When ACTE is removed, the MIoU scores drop slightly by approximately 1%, indicating that ACTE’s adaptive clustering contributes positively to the model’s accuracy by better aligning with data structures. The removal of TMA, however, a more pronounced decline of around 1.5%, underscoring the importance of mask-based attention in refining feature extraction. The results reveal that the ensemble of ACTE and TMA can provide the best performance, which indicates that the ACTE and TMA are complementary to each other.
5.6 Parameter Analysis
Figure 7: MIoU changes as $p_{\mathrm{T}}$ varies from 10% to 50%.
Building on the explanation of the transferable masked attention mechanism, the parameter $\lambda_{\mathrm{T}}$ is crucial as it determines which regions are considered transferable. Specifically, $\lambda_{\mathrm{T}}$ is set as the value corresponding to the $p_{\mathrm{T}}$ percentile of the elements in the transferability map $\bm{T}$. This means that a certain percentage, $p_{\mathrm{T}}$, of the regions with the higher transferability values are excluded from the attention mechanism. As shown in Fig. 7, when the parameter $p_{\mathrm{T}}$ gradually increases within the range of 10% to 30%, the model performance across all target domains exhibits a stable upward trend when using Cityscapes as the source domain, reaching its peak at $p_{\mathrm{T}}=30\%$. However, when $p_{\mathrm{T}}$ exceeds the critical value of 30%, the model performance undergoes a systematic decline. Notably, the model with GTAV as the target domain demonstrates a response pattern distinct from other models. Although its performance degradation is delayed by approximately 10% beyond $p_{\mathrm{T}}=30\%$, subsequent experimental results indicate that as $p_{\mathrm{T}}$ continues to increase to 50%, the MIoU drops by approximately 0.5% from its peak. Furthermore, throughout the high-parameter range , the model fails to surpass the previously achieved performance peak. Therefore, we ultimately choose 30% as the final value for $p_{\mathrm{T}}$.
6 Conclusion and Future work
This paper presents a novel framework for semantic segmentation transfer learning, leveraging the Adaptive Cluster-based Transferability Estimator (ACTE) and Transferable Masked Attention (TMA) to address the challenges of domain adaptation in vision transformers. ACTE dynamically evaluates region-level transferability, enabling targeted adaptation by focusing on the most informative and domain-divergent regions. TMA integrates transferability maps into Mask2Former’s attention mechanism, enhancing the model’s ability to prioritize regions with low transferability and semantic uncertainty. Extensive experiments across multiple benchmarks demonstrate the effectiveness of our approach, achieving significant improvements over both vanilla fine-tuning and state-of-the-art methods. Looking ahead, future work could investigate cross-modal adaptability and integrate emerging architectural innovations to further advance the field. These directions hold the potential to develop more intelligent and adaptable visual systems, capable of handling the complexities of real-world scenarios with greater efficiency and accuracy.